hartnn (hartnn):

if in certain language, we can write SKIN as 7(4)5(2)4(3)6(2) then what can we write TENT as ??

hartnn (hartnn):

@mathslover

OpenStudy (allank):

Puzzles...nice. Thinking...

hartnn (hartnn):

@shubhamsrg @ganeshie8

hartnn (hartnn):

@UnkleRhaukus @sauravshakya @Algebraic!

OpenStudy (unklerhaukus):

_ _ 6(2) _ ?

hartnn (hartnn):

yup.

OpenStudy (unklerhaukus):

i dont know

Parth (parthkohli):

Is this puzzle related to A = 1, B = 2.....Z = 26?

OpenStudy (allank):

I think I've figured it out:

hartnn (hartnn):

then message me the logic / answer, not here, let others figure it out..

OpenStudy (allank):

Okay.

mathslover (mathslover):

it's easy one ... wait

hartnn (hartnn):

yup, its easier .... u too msg me.

mathslover (mathslover):

I will message you if I will get it hartnn ...

mathslover (mathslover):

did allank message you? is he right?

OpenStudy (allank):

So, did mathsolver get it?

mathslover (mathslover):

wait

hartnn (hartnn):

@allank is very close!

Parth (parthkohli):

Math slover is a genius :)

mathslover (mathslover):

oh ...

mathslover (mathslover):

hint please

OpenStudy (allank):

Nice one @hartnn Do give @mathslover what I got, then he can help me figure out the rest of it

mathslover (mathslover):

yes please message me hartnn may be I will be able to ..

hartnn (hartnn):

jonask was also much close....but even he didn't get it exact

hartnn (hartnn):

HINT : 1(1)= . 2(1) = A

OpenStudy (allank):

Aha!

hartnn (hartnn):

@ash2326 can try this.

OpenStudy (shubhamsrg):

am clueless!! hmm..

mathslover (mathslover):

i am curious :P

hartnn (hartnn):

ok, MAYBE some of u might not be seeing or using this every day, but most of u , yes..

OpenStudy (anonymous):

Is it keys on mobile phone

OpenStudy (shubhamsrg):

ohh..hahahahaha...amazing !

OpenStudy (anonymous):

thats it

hartnn (hartnn):

yes!!

mathslover (mathslover):

:P good work saurav

OpenStudy (anonymous):

TENT=8(1)3(2)6(2)8(1)

OpenStudy (anonymous):

Thanx

mathslover (mathslover):

right

OpenStudy (allank):

Bravo, @sauravshakya

mathslover (mathslover):

T = (comes first ) = 8(1) E = (comes second) = 3(2) N = comes second = 6(2) T = again ..

OpenStudy (unklerhaukus):

great work, @sauravshakya!

OpenStudy (ash2326):

@sauravshakya awesome :)

OpenStudy (allank):

Somebody medal @hartnn too. Awesome question. :)

OpenStudy (anonymous):

THANX GUYS

hartnn (hartnn):

lol!

mathslover (mathslover):

1) qwerty keyboard 2) mobile keyboard 3) now which one hartnn :P

hartnn (hartnn):

whatever i see around!

ganeshie8 (ganeshie8):

hartnn is text addicted ;p

OpenStudy (allank):

True @ganeshie8 :) nobody just sits down and writes such a brain-pickler :)

hartnn (hartnn):

lol! i haven't sent a text in months!